# UniSite: The First Cross-Structure Dataset and Learning Framework for End-to-End Ligand Binding Site Detection

### v1

@@ 5.1 Settings
UniSite-DS is used for training and validation. We employ MMSeq2 (mmseqs2018, ) to ensure that no test UniProt sequence has similarity above 0.9 to any sequence in the training set. For each UniProt sequence, we select the PDB structure with the highest sequence identity as the representative structure. Additionally, we compare UniSite with baseline methods on widely-used binding site benchmark datasets, HOLO4K (p2rank2018, ) and COACH420 (p2rank2018, ). Following DeepSurf (deepsurf2021, ) and EquiPocket (equipocket2023, ), we use the mlig subsets of HOLO4K and COACH420 for evaluation. Since our models are trained under a UniProt-centric schema, we only consider single-chain structures, denoting the test datasets as HOLO4K-sc and COACH420 (all structures in COACH420 are originally single-chain). All test UniProt entries are strictly excluded from the training set. More details are provided in Appendix D.
@@
$\mathcal{L}_{\text{mask}}=\lambda_{\text{bce}}\mathcal{L}_{\text{bce}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}},$
@@
Table 1: Results on UniSite-DS. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket-rescore denotes sites initially predicted by Fpocket and subsequently rescored by P2Rank. ${}^{\text{b}}$ VN-EGNN only outputs centers of predicted sites. For each center, We include the residues within a 9Å radius, which has the best AP performance (Appendix E).
@@
For comprehensive evaluation, we compare IoU-based AP and traditional DCC/DCA metrics in HOLO4K-sc and COACH420. Since a Uniprot-centric data entry can contain ligands from multiply PDB structures, coordinate-dependent metrics like DCC and DCA become unsuitable due to potential inconsistencies in ligand spatial arrangements. Consequently, we merely use IoU-based AP metric on UniSite-DS. Following EquiPocket (equipocket2023, ), we set the DCC/DCA threshold to 4Å, and compute the DCC/DCA success rate of top-$\bm{n}$ predictions. We calculate AP using IoU thresholds of $0.3$ and $0.5$.
@@ 5.2 Results on UniSite-DS
The results on UniSite-DS are shown in Table 1. The geometry-based method Fpocket (fpocket2009, ) exhibits inferior performance since it merely considers the geometry and electronegativity. P2Rank (p2rank2018, ) achieves better results by extracting the protein surface features with Random Forest. DeepPocket (deeppocket2022, ) utilizes 3D-CNN to rescore and refine Fpocket predictions, improving the preformance in AP0.5. Notably, Fpocket-rescore, which combines Fpocket’s initial predictions with P2Rank’s re-ranking, surpasses both P2Rank and DeepPocket, highlighting the importance of proper scoring in binding site detection as captured by our AP metric. For graph models, GrASP (grasp2024, ) achieves further improvements in AP0.5 by employing graph attention networks. However, VN-EGNN (vnegnn2024, ) performs poorly under AP metrics, because it only outputs predicted binding site centers, discarding structural properties (shape and size) or residue identification. For evaluation purposes, we include the residues within a 9Å radius of each predicted center, which has the best AP performance (Appendix E).
@@ 5.3 Results on HOLO4K-sc and COACH420
Table 2: Results on HOLO4K-sc and COACH420. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket-rescore denotes sites initially predicted by Fpocket and subsequently rescored by P2Rank. ${}^{\text{b}}$ VN-EGNN only outputs centers of predicted sites. For each center, We include the residues within a 9Å radius, which has the best AP performance (Appendix E).
@@ 5.4 Ablation study
6 Conclusion
@@ Appendix A Impact of Binding Site Accuracy on Downstream Tasks
Appendix B Case Study of Classical Methods
@@ Appendix C Related Work
In recent years, deep learning methods have emerged for protein–ligand binding site detection. CNN-based approaches [12, 36, 40] treat protein structures as 3D images, applying 3D convolutional neural networks similar to those used in computer vision. Alternatively, GNN-based methods [15, 14, 13] utilize graph neural networks by constructing graphs incorporating both geometric and chemical features of proteins. Despite their improved performance, these methods typically adopt discontinuous workflows: they first perform semantic segmentation to generate binary masks of potential binding residues/atoms, then cluster these masks into discrete binding sites. This fragmented pipeline heavily depends on post-processing (e.g., clustering algorithms), inherently limiting end-to-end optimization and struggle with overlapping binding sites.
Appendix D External Datasets
@@
Appendix E The AP Evaluation of VN-EGNN
@@
Table S1: AP evaluation results about VN-EGNN with different radius.
@@
Appendix F Full Results on HOLO4K-sc and COACH420
Table S2: Full results on HOLO4K-sc. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket sites rescored by P2Rank. ${}^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix E).
@@ Appendix F Full Results on HOLO4K-sc and COACH420
Table S3: Full results on COACH420. We highlight the top two performing methods for each metric in bold. ${}^{\text{a}}$ Fpocket sites rescored by P2Rank. ${}^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix E).

### v2

@@ 5.1 Settings
UniSite-DS is used for training and validation. We employ MMSeq2 (mmseqs2018, ) to ensure that no test UniProt sequence has similarity above 0.9 to any sequence in the training set. For each UniProt sequence, we select the PDB structure with the highest sequence identity as the representative structure. Additionally, we compare UniSite with baseline methods on widely-used binding site benchmark datasets, HOLO4K (p2rank2018, ) and COACH420 (p2rank2018, ). Following DeepSurf (deepsurf2021, ) and EquiPocket (equipocket2023, ), we use the mlig subsets of HOLO4K and COACH420 for evaluation. Since our models are trained under a UniProt-centric schema, we only consider single-chain structures, denoting the test datasets as HOLO4K-sc and COACH420 (all structures in COACH420 are originally single-chain). All test UniProt entries are strictly excluded from the training set. More details are provided in Appendix F.
@@
$\mathcal{L}_{\text{mask}}=\lambda_{\text{bce}}\mathcal{L}_{\text{bce}}+\lambda_{\text{dice}}\mathcal{L}_{\text{dice}}$ (8)
@@
Table 1: Results on UniSite-DS. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket-rescore denotes sites initially predicted by Fpocket and subsequently rescored by P2Rank. $^{\text{b}}$ VN-EGNN only outputs centers of predicted sites. For each center, We include the residues within a 9Å radius, which has the best AP performance (Appendix H).
@@
For comprehensive evaluation, we compare IoU-based AP and traditional DCC or DCA metrics in HOLO4K-sc and COACH420. Since a Uniprot-centric data entry can contain ligands from multiply PDB structures, coordinate-dependent metrics like DCC and DCA become unsuitable due to potential inconsistencies in ligand spatial arrangements. Consequently, we merely use IoU-based AP metric on UniSite-DS. Following EquiPocket (equipocket2023, ), we set the DCC or DCA threshold to 4Å, and compute the DCC or DCA success rate of top-$\bm{n}$ predictions. We calculate AP using IoU thresholds of $0.3$ and $0.5$.
@@ 5.2 Results on UniSite-DS
The results on UniSite-DS are shown in Table 1. The geometry-based method Fpocket (fpocket2009, ) exhibits inferior performance since it merely considers the geometry and electronegativity. P2Rank (p2rank2018, ) achieves better results by extracting the protein surface features with Random Forest. DeepPocket (deeppocket2022, ) utilizes 3D-CNN to rescore and refine Fpocket predictions, improving the preformance in AP0.5. Notably, Fpocket-rescore, which combines Fpocket’s initial predictions with P2Rank’s re-ranking, surpasses both P2Rank and DeepPocket, highlighting the importance of proper scoring in binding site detection as captured by our AP metric. For graph models, GrASP (grasp2024, ) achieves further improvements in AP0.5 by employing graph attention networks. However, VN-EGNN (vnegnn2024, ) performs poorly under AP metrics, because it only outputs predicted binding site centers, discarding structural properties (shape and size) or residue identification. For evaluation purposes, we include the residues within a 9Å radius of each predicted center, which has the best AP performance (Appendix H).
@@ 5.3 Results on HOLO4K-sc and COACH420
Table 2: Results on HOLO4K-sc and COACH420. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket-rescore denotes sites initially predicted by Fpocket and subsequently rescored by P2Rank. $^{\text{b}}$ VN-EGNN only outputs centers of predicted sites. For each center, We include the residues within a 9Å radius, which has the best AP performance (Appendix H).
@@ 5.4 Ablation study
6 Concluding Remarks and Future Perspectives
Key Contributions.
@@ 6 Concluding Remarks and Future Perspectives
Limitations and Future Work.
The current version of UniSite-DS involves manual curation to remove unreasonable entries. A promising direction for future work is to develop automated methods for repairing and reintegrating excluded data to further enhance the dataset’s coverage and quality. Additionally, our current model design aims to demonstrate the effectiveness of the end-to-end ligand binding site learning framework, without incorporating specialized feature engineering. Future investigations could explore the inclusion of specialized feature engineering to further improve model performance and generalization ability.
Acknowledgements
We thank the reviewers for their constructive feedback, and Hannes Stärk, Chenyu Wang, Zuobai Zhang, Zaixi Zhang, and Bozitao Zhong for helpful discussions. Liwei Wang is supported by National Science and Technology Major Project (2022ZD0114902) and National Science Foundation of China (NSFC92470123, NSFC62276005).
@@ Appendix A Impact of Binding Site Accuracy on Downstream Tasks
Appendix B UniSite-DS Curation Workflow and Representative Manual Inspection Case Studies
As shown in Figure S2, the UniSite-DS workflow comprises three main components: (I) dataset curation, (II) quality control, and (III) manual inspection. Below, we describe two representative types of cases encountered during manual inspection that required special consideration:
  1. 1.
    Case 1. Supramolecular assemblies across multiple protein subunits. The Photosystem I-LHCI Supercomplex is a core component of the photosynthetic machinery, consisting of multiple subunits such as UniProt ID: P05310 (Photosystem I P700 chlorophyll a apoprotein A1, PDB: 7dkz_A), Q41038 (Chlorophyll a-b binding protein, 7dkz_B), and Q32904 (Chlorophyll a-b binding protein 3, 7dkz_C). Although each of these proteins binds a large number of Chlorophyll a molecules, they function as part of a tightly integrated supramolecular assembly. Therefore, they are not suitable to be included as independent entries in the UniSite-DS database and have been excluded in the current version.
  2. 2.
    Case 2. Large composite cavities jointly formed by multiple ligands. Trypanothione reductase (UniProt ID: Q389T8) is a key enzyme in Trypanosoma brucei that specifically catalyzes the reduction of trypanothione, functionally analogous to glutathione reductase in mammals. The ligands from structures PDB: 5s9x_A, 5s9t_A, and 2wov_C together form a large binding cavity. These ligand binding sites could potentially be merged. However, since each ligand actually occupies only a portion of the cavity rather than the entire cavity, whether such a composite cavity formed by different ligands should be considered a unified binding site often depends on system-specific definitions in the literature. As a result, entries that require further literature-based validation were excluded from the current version of the UniSite-DS database.
Figure S2: Overview of the UniSite-DS workflow. Workflow of (I) dataset curation, (II) quality control, and (III) manual inspection for UniSite-DS, together with (IV) representative manual inspection case studies.
@@ Appendix C Related Work
In recent years, deep learning methods have emerged for protein–ligand binding site detection. CNN-based approaches [12, 37, 41] treat protein structures as 3D images, applying 3D convolutional neural networks similar to those used in computer vision. Alternatively, GNN-based methods [15, 14, 13] utilize graph neural networks by constructing graphs incorporating both geometric and chemical features of proteins. Despite their improved performance, these methods typically adopt discontinuous workflows: they first perform semantic segmentation to generate binary masks of potential binding residues/atoms, then cluster these masks into discrete binding sites. This fragmented pipeline heavily depends on post-processing (e.g., clustering algorithms), inherently limiting end-to-end optimization and struggling with overlapping binding sites. DETR-based architectures have also been adapted for other protein tasks, such as ProtDETR [59], which performs enzyme function classification rather than binding site detection.
Appendix D Baseline Justification
We selected representative baseline methods that span the major methodological paradigms in ligand binding site detection. Fpocket [8] represents traditional geometry-based computational algorithms. P2Rank [11] is the most widely used machine learning approach. Among deep learning approaches, DeepPocket [12], GrASP [14], and VN-EGNN [15] were included as they represent different neural network architectures: DeepPocket employs 3D CNNs, while GrASP and VN-EGNN utilize GNNs.
For a fair and consistent comparison, we extracted ligand binding site residues from each baseline’s output files according to their respective standard formats:
  • Fpocket: For each predicted pocket, Fpocket outputs a file named pocket{index}_atm.pdb, which contains the atomic coordinates of the predicted binding site in PDB format. These atoms were directly used to identify the corresponding binding residues.
  • P2Rank: P2Rank generates a {name}.pdb_prediction.csv file for each input structure. The residue_ids column specifies the chain IDs and residue IDs of residues constituting each predicted binding site. These identifiers were parsed to obtain the binding site residues.
  • DeepPocket: DeepPocket refines and re-scores the pockets predicted by Fpocket, while maintaining the same output format.
  • GrASP: GrASP outputs a {name}_probs.pdb file, where the bfactor column encodes the predicted binding probabilities of heavy atoms. Following the original publication, we filtered atoms based on their predicted scores and clustered them spatially into distinct binding sites.
  • VN-EGNN: VN-EGNN produces a prediction.csv file containing four columns (x, y, z, rank) that specify the 3D coordinates and ranking of predicted pocket centers. As the method does not directly output residue-level predictions, binding site residues were obtained by including all residues within a defined radius, as described in Appendix H.
Appendix E Case Study of Classical Methods
Appendix F External Datasets
@@
Appendix G Quantitative Analysis of Evaluation Flaws in DCC and DCA Metrics
In Section 4, we discussed two critical limitations in the design of the DCC and DCA metrics. Here, we conduct a quantitative analysis on the HOLO4K-sc benchmark to demonstrate the flawed evaluation introduced by these metrics.
Limitation 1. The absence of proper matching criteria may lead to double-counting of predictions (Figure 4 A). We quantified the proportion of proteins affected by double-counting during evaluation on the HOLO4K-sc benchmark (Table S1). The results reveal that DCC and DCA metrics suffer from widespread double counting artifacts, which significantly distort model performance assessment.
Table S1: Double counting (DC) rate of DCC or DCA metrics on HOLO4K-sc. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket sites rescored by P2Rank. $^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix H).
Method HOLO4K-sc
DC of $\text{DCC}_{\text{top-}n}$ DC of $\text{DCA}_{\text{top-}n}$
Fpocket [8] 18.80% 18.31%
[…]
P2Rank [11] 18.98% 18.31%
DeepPocket [12] 13.53% 14.21%
GrASP [14] 16.29% 14.88%
[…]
UniSite-1D (ours) 8.88% 8.94%
UniSite-3D (ours) 9.86% 10.04%
Limitation 2. DCC or DCA only evaluate the center of binding sites and are ligand-dependent, which leads to evaluation failures in certain scenarios (Figure 4 B-D). For HOLO4K-sc, we calculated these metrics of the centroid of ground truth binding residues for each protein. The results indicate that the mean ground truth DCC is 2.15 Å (92.65% < 4 Å), and the mean ground truth DCA is 1.57 Å (98.88% < 4 Å). However, in principle, both DCC and DCA should ideally be 0 when evaluated using ground truth binding residues, indicating these metrics inherently contain systematic bias.
To mitigate some of DCC’s inherent limitations, we defined a corrected metric, DCC-residue, which uses the center of ground truth binding residues rather than the ligand center for calculation. This modification resolves failure cases caused by ligand diversity in traditional DCC evaluation. As shown in Table S2, the corrected DCC-residue metric exhibits improved consistency with IoU-based AP in ranking the performance of different methods.
Table S2: IoU-based AP, DCC-residue and DCC results on HOLO4K-sc. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket sites rescored by P2Rank. $^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix H).
Method HOLO4K-sc
$\text{AP}_{0.3}\uparrow$ $\text{DCC-residue}_{\text{top-}n}\uparrow$ $\text{DCC}_{\text{top-}n}\uparrow$
Fpocket [8] 0.2711 0.2982 0.3076
[…]
P2Rank [11] 0.6011 0.4972 0.5300
DeepPocket [12] 0.5415 0.4902 0.4925
GrASP [14] 0.6668 0.5379 0.5131
[…]
UniSite-1D (ours) 0.6867 0.6400 0.5538
UniSite-3D (ours) 0.7091 0.6264 0.5716
Appendix H The AP Evaluation of VN-EGNN
@@
Table S3: AP evaluation results of VN-EGNN using different radii.
@@
Appendix I The Recall Evaluation of Different Methods
The AP metric provides a comprehensive evaluation of binding site detection performance by considering both precision and recall. However, it is also important to evaluate how many true binding sites can be recovered under different IoU thresholds, since the goal of binding site detection is to identify novel and biologically meaningful sites/pockets. We computed the Recall of different methods across multiple IoU thresholds on the HOLO4K-sc dataset, and the results are summarized in Table S4. UniSite-3D achieves state-of-the-art Recall across all IoU thresholds.
Notably, the high recall obtained by Fpocket is due to its tendency to output nearly all potential cavities in a protein, rather than accurately identifying biologically relevant pockets. In fact, Fpocket tends to assign lower scores to true binding sites, making it challenging for biologists to distinguish meaningful pockets. This limitation has motivated many studies to develop rescoring strategies for Fpocket. While Recall reflects the potential of a method to detect biologically significant pockets, it does not account for the confidence score of each prediction. This limitation is addressed by the AP metric, which is why we adopt it as a fair and balanced criterion for evaluating different methods.
Table S4: The AP and Recall results on HOLO4K-sc. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket sites rescored by P2Rank. $^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix H).
Method HOLO4K-sc
$\text{AP}_{0.3}$ $\text{AP}_{0.5}$ $\text{Recall}_{0.3}$ $\text{Recall}_{0.5}$ $\text{Recall}_{0.7}$ $\text{Recall}_{0.9}$
Fpocket [8] 0.2711 0.1488 0.8361 0.5922 0.2130 0.0253
[…]
P2Rank [11] 0.6011 0.2625 0.7814 0.5337 0.1868 0.0089
DeepPocket [12] 0.5415 0.2891 0.7514 0.5824 0.2584 0.022
GrASP [14] 0.6668 0.4126 0.7186 0.5374 0.2537 0.0159
[…]
UniSite-1D (ours) 0.6867 0.4595 0.8212 0.6199 0.3535 0.0824
UniSite-3D (ours) 0.7091 0.5446 0.8469 0.6901 0.4106 0.1039
Appendix J Full Results on HOLO4K-sc and COACH420
Table S5: Full results on HOLO4K-sc. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket sites rescored by P2Rank. $^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix H).
@@ Appendix I The Recall Evaluation of Different Methods
Table S6: Full results on COACH420. We highlight the top two performing methods for each metric in bold. $^{\text{a}}$ Fpocket sites rescored by P2Rank. $^{\text{b}}$ Residues within 9Å of each VN-EGNN predicted center, which has the best AP performance (Appendix H).
@@
Appendix K Pseudo-code for Average Precision Calculation
Algorithm 1 Average Precision Calculation
0:Input:
0:   prediction_list: A list, where each element represents the predictions for one protein. Each element is of the form $\{(s_{i},m_{i})|s_{i}\in\mathbb{R},m_{i}\in\{0,1\}^{L}\}_{i=1}^{N}$, where $m_{i}$ represents the $i$-th predicted binding site as a binary mask of length $L$, and $s_{i}$ denotes the $i$-th confidence score.
0:   ground_truth_list: A list, where each element represents the ground truth (gt) binding sites for one protein. Each element is of the form $\{m_{i}^{gt}|m_{i}^{gt}\in\{0,1\}^{L}\}_{i=1}^{N_{gt}}$, where $m_{i}^{gt}$ represents the $i$-th ground truth binding site as a binary mask of length $L$.
0:   iou_threshold: IoU threshold for considering a prediction as True Positive.
0:Output: AP: Average Precision under the given IoU threshold. # Step 1: Determine TP (True Positive) or FP (False Positive) for each prediction
1:for predictions_per_protein, ground_truths_per_protein in ZIP(prediction_list, ground_truth_list) do
2:  Set all ground truths in ground_truths_per_protein as unused
[…]
4:  for $i=1$ to NN do
5:   Find ground_truth_j that has the max residue-level IoU($m_{j}^{gt}$, $m_{i}$) with prediction_i
6:   if IoU($m_{j}^{gt}$, $m_{i}$) >> iou_threshold and ground_truth_j is unused then
7:    Mark prediction_i as TP
8:    Set ground_truth_j as used
9:   else
10:    Mark prediction_i as FP
11:   end if
12:  end for
13:end for# Step 2: Sort all predictions across all proteins by decreasing confidence scores
14:all_predictions $\leftarrow$ FLATTEN_ALL(prediction_list)
15:all_ground_truths $\leftarrow$ FLATTEN_ALL(ground_truth_list)
16: Sort all_predictions by decreasing confidence scores # Step 3: Calculate the precision-recall curve
17: Set cum_TP, cum_FP as 0, precision_list and recall_list as empty
18:for prediction_i in all_predictions do
19:  if prediction_i is marked as TP then
20:   cum_TP $\leftarrow$ cum_TP + 1
21:  else
22:   cum_FP $\leftarrow$ cum_FP + 1
23:  end if
24:  precision_list.append(cum_TP / (cum_TP + cum_FP))
25:  recall_list.append(cum_TP / LEN(all_ground_truths))
26:end for# Step 4: Calculate average precision
27: Calculate AP as the area under the precision-recall curve
28:return AP